# Optimal and Fair Encouragement Policy Evaluation and Learning

### v4

@@ A.2 Additional related work
Intention-to-treat analysis. We appeal to intention-to-treat analysis with randomness that either arises from human decision-makers or individual non-adherence/non-compliance, but we generally assume the data does not include information about the identity of different decision-makers, which is common with publicly available data. Our conditional exclusion restriction also means that certain decomposed effects are zero, so mediation analysis is less relevant. A related literature studies principal stratification Jiang et al. [2020], which is less interpretable since stratum membership is unknown. Similarly, even though encouragement effects are driven by compliers, complier-conditional analysis is less policy-relevant since complier identities are unknown. In general, our causal identification arguments are based on covariate-adjusted intention-to-treat analysis and covariate-adjusted as-treated analysis. We avoid estimation of stratum-specific effects, because if complier status is unknown, prescriptive decision rules cannot directly personalize by stratum membership.
@@
Fair off-policy learning. We highlight some most closely related works in off-policy learning (omitting works in the sequential setting). [Metevier et al., 2019] studies high-probability fairness constraint satisfaction. [Kim et al., 2022] studies doubly-robust causal fair classification, while others have imposed deterministic resource constraints on the optimal policy formulation [Chohlas-Wood et al., 2021]. Other works study causal or counterfactual risk assessments [Mishler et al., 2021, Coston et al., 2020]. Our perspective is closer to that of off-policy learning, i.e. approximating direct control over the intervention by assuming stability in decision-maker treatment assignment probabilities. [Kallus and Zhou, 2019] studies (robust) bounds for treatment responders in binary outcome settings; this desiderata is coupled to classification notions of direct treatment. Again, our focus is on modeling the fairness implications of non-adherence. Indeed, in order to provide general algorithms and methods, we do build on prior fair classification literature. A different line of work studies “counterfactual" risk assessments which models a different concern.
Principal stratification and mediation analysis in causal inference
@@
[Liu et al., 2021] studies an optimal test-and-treat regime under a no-direct-effect assumption, that assigning a diagnostic test has no effect on outcomes except via propensity to treat, and studies semiparametric efficiency using Structural Nested-Mean Models. Though our exclusion restriction is also a no-direct-effect assumption, our optimal treatment regime is in the space of recommendations only as we do not have control over the final decision-maker, and we consider generally nonparametric models.
@@
Human-in-the-loop in consequential domains. There is a great deal of interest in designing algorithms for the “human in the loop" and studying expertise and discretion in human oversight in consequential domains [De-Arteaga et al., 2020]. On the algorithmic side, recent work focuses on frameworks for learning to defer or human-algorithm collaboration. Our focus is prior to the design of these procedures for improved human-algorithm collaboration: we primarily hold fixed current human responsiveness to algorithmic recommendations. Therefore, our method can be helpful for optimizing local nudges. Incorporating these algorithmic design ideas would be interesting directions for future work.
Empirical literature on judicial discretion in the pretrial setting.
Studying a slightly different substantive question, namely causal effects of pretrial decisions on later outcomes, a line of work uses individual judge decision-makers as a leniency instrumental variable for the treatment effect of (for example, EM) on pretrial outcomes [Arnold et al., 2022, 2018, Lum et al., 2017]. And, judge IVs rely on quasi-random assignment of individual judges. We focus on the prescriptive question of optimal recommendation rules in view of patterns of judicial discretion, rather than the descriptive question of causal impacts of detention on downstream outcomes.
@@
Appendix B Additional discussion on method
B.1 Extrapolation in ˜2
@@
Proposition 7 (Control variate for alternative identification ).
@@
$\textstyle V(\pi)=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}\left[\left\{c_{rt}(Y(t))\frac{\mathbb{I}\left[T=t\right]}{p_{t}(X)}+\left(1-\frac{\mathbb{I}\left[T=t\right]}{p_{t}(X)}\right)\mu_{t}(X)\right\}p_{t\mid r}(X)\right]$
@@
Appendix C Additional discussion on general optimization method
C.1 Additional discussion on constrained optimization
@@
$V_{\epsilon}^{*}=\max_{\pi}\;\{\mathbb{E}[c(\pi,T(\pi),Y(\pi))]\colon\mathbb{E}[T(\pi)\mid A=a]-\mathbb{E}[T(\pi)\mid A=b]\leq\epsilon\}$ (3)
C.2 Additional discussion on Algorithm˜2 (general algorithm)
C.2.1 Weighted classification reduction and off-policy estimation
@@
C.2.2 Additional fairness constraints and examples in this framework
In this section we discuss additional fairness constraints and how to formulate them in the generic framework. Much of this discussion is quite similar to [Agarwal et al., 2018] (including in notation) and is included in this appendix for completeness only. We only additionally provide novel identification results for another fairness measure on causal policies in Section˜C.2.3, concrete discussion of the reduction to weighted classification, and provide concrete descriptions of the causal fairness constraints in the more general framework.
@@ C.2.2 Additional fairness constraints and examples in this framework
Example 5 (Writing treatment parity in the general constrained classification framework.).
@@
We can write this in the generic notation given previously by letting $\mathcal{J}={\mathcal{A}\cup\{\circ\}},$
$g_{j}(O,\pi(X);\eta)=\pi_{1}(X)(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))+p_{1\mid 0}(X,A),\forall j.$
We let the conditioning events $\mathcal{E}_{a}=\{A=a\},\mathcal{E}_{\circ}=\{\text{True}\},$ i.e. conditioning on the latter is equivalent to evaluating the marginal expectation. Then we express Equation˜4 as a set of equality constraints $h_{a}(\pi)=h_{\circ}(\pi),$ leading to pairs of inequality constraints,
$\left\{\begin{matrix}h_{u}(\pi)-h_{\circ}(\pi)\leq 0\\ h_{\circ}(\pi)-h_{u}(\pi)\leq 0\end{matrix}\right\}_{u\in\mathcal{A}}$
The corresponding coefficients of $M$ over this enumeration over groups $(\mathcal{A})$ and epigraphical enforcement of equality $(\{+,-\})$ equation (1), gives $\mathcal{K}=\mathcal{A}\times\{+,-\}$ so that $M_{(a,+),a^{\prime}}=\mathbf{1}\left\{a^{\prime}=a\right\},M_{(a,+),\star}=-1$, $M_{(a,-),a^{\prime}}=-\mathbf{1}\left\{a^{\prime}=a\right\},M_{(a,-),\star}=1$, and $\mathbf{d}=\mathbf{0}$. Further we can relax equality to small amounts of constraint relaxation by instead setting $d_{k}>0$ for some (or all) $k$.
Next, we discuss a more complicated fairness measure. We first discuss identification and estimation before we also describe how to incorporate it in the generic framework.
C.2.3 Responder-dependent fairness measures
We consider a responder framework on outcomes (under our conditional exclusion restriction). Because the contribution to the AEE is indeed from the responder strata, this corresponds to additional estimation of the responder stratum.
We enumerate the four possible realizations of potential outcomes (given any fixed recommendation) as $(Y(0(r)),Y(1(r))\in$ $\{0,1\}^{2}$. We call units with $(Y(0(r)),Y(1(r)))=(0,1)$ responders, $(Y(0(r)),Y(1(r)))=(1,0)$ anti-responders, and $Y(0(r))=Y(1(r))$ non-responders. Such a decomposition is general for the binary setting.
Assumption 9 (Binary outcomes, treatment).
$T,Y\in\{0,1\}$
Assumption 10 (Monotonicity).
$Y(T(1))\geq Y(T(0))$
Importantly, the conditional exclusion restriction of Assumption˜2 implies that responder status is independent of recommendation. Conditional on observables, whether a particular individual is a responder is independent of whether someone decides to treat them when recommended. In this way, we study responder status analogous to its use elsewhere in disparity assessment in algorithmic fairness [Imai et al., 2020, Kallus et al., 2019a]. Importantly, this assumption implies that the conditioning event (of being a responder) is therefore independent of the policy $\pi$, so it can be handled in the same framework. s
We may consider reducing disparities in resource expenditure given responder status.
We may be interested in the probability of receiving treatment assignment given responder status.
Example 6 (Fair treatment expenditure given responder status).
$\mathbb{E}[T(\pi)\mid Y(1(R))>Y(0(R)),A=a]-\mathbb{E}[T(\pi)\mid Y(1(R))>Y(0(R)),A=b]\leq\epsilon$
We can obtain identification via regression adjustment:
Proposition 8 (Identification of treatment expenditure given responder status).
Assume Assumptions˜9 and 10.
$P(T(\pi)=1\mid A=a,Y(1(\pi))>Y(0(\pi)))=\frac{\sum_{r}\mathbb{E}[\pi_{r}(X)p_{1\mid r}(X,A)(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=a])}{\mathbb{E}[(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=a]}$
Therefore this can be expressed in the general framework.
Example 7 (Writing treatment responder-conditional parity in the general constrained classification framework.).
For each $u\in\mathcal{A}$ we enforce that
$\frac{\sum_{r}\mathbb{E}[\pi_{r}(X)p_{1\mid r}(X,A)(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=u])}{\mathbb{E}[(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=u]}=\frac{\sum_{r}\mathbb{E}[\pi_{r}(X)p_{1\mid r}(X,A)(\mu_{1}(X,A)-\mu_{0}(X,A))])}{\mathbb{E}[(\mu_{1}(X,A)-\mu_{0}(X,A))]}$
We can write this in the generic notation given previously by letting $\mathcal{J}={\mathcal{A}\cup\{\circ\}},$
$g_{j}(O,\pi(X);\eta)=\frac{\{\pi_{1}(X)(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))+p_{1\mid 0}(X,A)\}(\mu_{1}(X,A)-\mu_{0}(X,A))}{\mathbb{E}[(\mu_{1}(X,A)-\mu_{0}(X,A))\mid A=a]},\forall j.$
Let $\mathcal{E}_{a}^{j}=\{A=a_{j}\},\mathcal{E}_{\circ}=\{\text{True}\},$ and we express Equation˜4 as a set of equality constraints of the above moment $h_{a}(\pi)=h_{\circ}(\pi),$ leading to pairs of inequality constraints,
$\left\{\begin{matrix}h_{u}(\pi)-h_{\circ}(\pi)\leq 0\\ h_{\circ}(\pi)-h_{u}(\pi)\leq 0\end{matrix}\right\}_{u\in\mathcal{A}}$
The corresponding coefficients of $M$ proceed analogously as for treatment parity.
C.2.4 Best-response oracles

### v5

@@ 8.1 Additional Related work
Intention-to-treat analysis. We appeal to intention-to-treat analysis with randomness that either arises from human decision-makers or individual nonadherence/noncompliance, but we generally assume the data does not include information about the identity of different decision-makers, which is common with publicly available data. Our conditional exclusion restriction also means that certain decomposed effects are zero, so mediation analysis is less relevant. A related literature studies principal stratification [Jiang et al., 2020], which is less interpretable since stratum membership is unknown. Similarly, even though encouragement effects are driven by compliers, complier-conditional analysis is less policy-relevant since complier identities are unknown. In general, our causal identification arguments are based on covariate-adjusted intention-to-treat analysis and covariate-adjusted as-treated analysis. We avoid estimation of stratum-specific effects, because if complier status is unknown, prescriptive decision rules cannot directly personalize by stratum membership.
@@
Fair off-policy learning. We highlight some most closely related works in off-policy learning (omitting works in the sequential setting). [Metevier et al., 2019] studies high-probability fairness constraint satisfaction. [Kim et al., 2022] studies doubly-robust causal fair classification, while others have imposed deterministic resource constraints on the optimal policy formulation [Chohlas-Wood et al., 2021]. Other works study causal or counterfactual risk assessments [Mishler et al., 2021, Coston et al., 2020]. Our perspective is closer to that of off-policy learning, i.e. approximating direct control over the intervention by assuming stability in decision-maker treatment assignment probabilities. [Kallus and Zhou, 2019] studies (robust) bounds for treatment responders in binary outcome settings; this desiderata is coupled to classification notions of direct treatment. Again, our focus is on modeling the fairness implications of non-adherence. Indeed, in order to provide general algorithms and methods, we do build on prior fair classification literature. A different line of work studies “counterfactual” risk assessments which models a different concern.
@@
Principal stratification and mediation analysis in causal inference. [Liu et al., 2021] studies an optimal test-and-treat regime under a no-direct-effect assumption, that assigning a diagnostic test has no effect on outcomes except via propensity to treat, and studies semiparametric efficiency using Structural Nested-Mean Models. Though our exclusion restriction is also a no-direct-effect assumption, our optimal treatment regime is in the space of recommendations only as we do not have control over the final decision-maker, and we consider generally nonparametric models.
@@
Other work on algorithmic fairness encouraging access in operations. There is extensive work on algorithmic fairness in operations, measured in different ways. bertsimas studies efficiency trade-offs versus Nash social welfare fairness. On the non-algorithmic advice, but encouragement side, Freund and Hssaine [2025] consider dynamic resource allocation under fairness constraints, when deciding monetary incentives for retention under stochastic participation. Other models of stochastic resource usage and demand are also relevant, though potentially more specialized than our non-adherence setting.
Human-in-the-loop in consequential domains. There is a great deal of interest in designing algorithms for the “human in the loop” and studying expertise and discretion in human oversight in consequential domains [De-Arteaga et al., 2020]. On the algorithmic side, recent work focuses on frameworks for learning to defer or human-algorithm collaboration. Our focus is prior to the design of these procedures for improved human-algorithm collaboration: we primarily hold fixed current human responsiveness to algorithmic recommendations. Therefore, our method can be helpful for optimizing local nudges. Incorporating these algorithmic design ideas would be interesting directions for future work.
Empirical literature on judicial discretion in the pretrial setting. Studying a slightly different substantive question, namely causal effects of pretrial decisions on later outcomes, a line of work uses individual judge decision-makers as a leniency IV for the treatment effect of (for example, EM) on pretrial outcomes [Arnold et al., 2022, 2018, Lum et al., 2017]. And, judge IVs rely on quasi-random assignment of individual judges. We focus on the prescriptive question of optimal recommendation rules in view of patterns of judicial discretion, rather than the descriptive question of causal impacts of detention on downstream outcomes.
@@
8.2 Alternative identification
Alternative IV-style interpretation following Wang and Tchetgen Tchetgen.
If one prefers to avoid our stronger treatment-unconfoundedness assumption and the corresponding heterogeneous-treatment-effect language, an alternative route available when the recommendation is truly randomized is to interpret the recommendation $R$ as an instrumental variable for treatment $T$. In our notation, let
$\Delta_{Y}(x,a):=\mathbb{E}[Y\mid R=1,X=x,A=a]-\mathbb{E}[Y\mid R=0,X=x,A=a]$
The corresponding conditional Wald estimand is
$\delta_{W}(x,a):=\frac{\Delta_{Y}(x,a)}{\kappa(x,a)}.$
[…]
$\mathbb{E}[T\mid R=1,X,A,U]-\mathbb{E}[T\mid R=0,X,A,U]=\mathbb{E}[T\mid R=1,X,A]-\mathbb{E}[T\mid R=0,X,A],$
that is, no additive UURR interaction in the first stage, or
$\mathbb{E}[Y(1)-Y(0)\mid X,A,U]=\mathbb{E}[Y(1)-Y(0)\mid X,A],$
that is, no additive UUtt interaction in the treatment effect. Under these assumptions,
$\delta_{W}(x,a)=\mathbb{E}[Y(1)-Y(0)\mid X=x,A=a].$
Interpretation relative to our main analysis.
Under our main text assumptions, we write the encouragement score as
$s(x,a)=\Delta_{Y}(x,a)=\kappa(x,a)\tau(x,a),$
where $\tau(x,a)$ is an as-treated heterogeneous treatment effect identified by treatment unconfoundedness. Under the Wang–Tchetgen Tchetgen alternative, the same reduced-form score is instead written as
$s(x,a)=\Delta_{Y}(x,a)=\kappa(x,a)\delta_{W}(x,a).$
Thus the policy-relevant reduced-form score $s(x,a)$ and the first-stage term $\kappa(x,a)$ are unchanged; what changes is the interpretation of the multiplier. In particular, one replaces the language of “heterogeneous treatment effects” with the language of conditional Wald effects. If one further strengthens the IV assumptions to a causal IV assumption and adds monotonicity,
[…]
then the same $\delta_{W}(x,a)$ also admits a conditional LATE (CLATE) interpretation, so that one may equivalently write $s(x,a)=\kappa(x,a)\tau_{c}(x,a),$ with $\tau_{c}(x,a)$ denoting a conditional complier-local average treatment effect.Then
$\tau_{c}(x,a):=\mathbb{E}\!\left[Y(1)-Y(0)\mid T(1)>T(0),\,X=x,\,A=a\right].$
Thus, under monotonicity, the term multiplying the first stage is no longer an as-treated heterogeneous treatment effect for the full stratum $(x,a)$, but rather the average treatment effect for the latent subgroup of compliers within that stratum. In this case, $\kappa(x,a)=\Pr(T(1)>T(0)\mid X=x,A=a)$ is the conditional complier mass, and $s(x,a)=\kappa(x,a)\tau_{c}(x,a)$ decomposes the value of encouragement into an extensive-margin term (how many are induced into treatment) and an intensive-margin term (how much treatment helps those induced units). Therefore the policy ranking is unchanged, but the interpretation of treatment-effect heterogeneity becomes local to compliers rather than population-wide.
Under all such alternative specifications, the core message remains the same: naive ITT targeting depends on a product of a term describing first-stage selection into treatment (whether they are restricted to compliers, as in the standard causal IV case with monotonicity, or not, without monotonicity) and a term describing treatment effectiveness (whether conditional to compliers or not).
9 Additional discussion on Section 4
9.1 A simple model with low fairness-value tradeoff.
The model
To concretize, consider a simple model with covariates partitioned as $(X^{s},X^{\kappa},X^{\tau})$ where $X$ is a shared factor affecting both compliance and heterogeneous treatment effect, $X^{\kappa}$ modifies compliance alone, and $X^{\tau}$ modifies the heterogeneous treatment effect alone. Consider the following model:
$X=(X^{s},X^{\tau},X^{\kappa}),\;\;\;X^{s}\sim N(0,1),\;\;\;X^{\tau}\sim N(0,1),\;\;\;X^{\kappa}\mid A=g\sim N(\mu_{g},1)$ (5)
$\tau(X,A)=\log(1+\exp(\beta_{s}X^{s}+\beta_{\tau}X^{\tau})),\qquad p_{1\mid 0}(X,A)=\bar{p}_{0},\quad\kappa(X,A)=(1-\bar{p}_{0})\Phi(\theta^{\kappa}X^{\kappa}+\theta^{s}X^{s})$ (6)
Importantly, in the simple model above, the protected attribute $A$ has no direct effect on compliance or the heterogeneous treatment effect. However, $A$ impacts the distribution of a compliance-relevant covariate $X^{\kappa}$. For example, $X^{\kappa}$ could be zip code that affects location and therefore distance to medical services. With other health information about a person, zip code wouldn’t plausibly biologically affect treatment effectiveness, but could affect adherence/compliance.
The simple model is primarily diagnostic: it shows that large disparities under naive ITT targeting can arise even when treatment benefits are identical across groups. As a secondary implication, it also identifies a regime in which suppressing compliance-only proxies can achieve a favorable fairness–value tradeoff, thereby motivating the fairness-constrained methods developed later. However, it may not be possible to implement group-specific thresholds and it may be procedurally fairer and preferable to optimize policies over $x$ alone rather than $x,a$. Let $\pi_{DP}^{blind}(X)$ denote a solution to eq. 3 but optimizing over policies that are measurable with respect to $(X^{s},X^{\tau})$.
Proposition 9.1
Under the above model of Equation 5,
$\mathbb{E}[\kappa(X,A)\mid X^{s},A=g]=(1-\bar{p}_{0})\Phi\left(\frac{\theta^{\kappa}\mu_{g}+\theta^{s}X^{s}}{\sqrt{1+\left(\theta^{\kappa}\right)^{2}}}\right),\;\;\mathbb{E}[\kappa(X,A)\mid A=g]=(1-\bar{p}_{0})\Phi\left(\frac{\theta^{\kappa}\mu_{g}}{\sqrt{1+\left(\theta^{\kappa}\right)^{2}+\left(\theta^{s}\right)^{2}}}\right)$
We also have that $\mathbb{E}[\pi_{DP}^{blind}(X^{s},X^{\tau})\mid A=a]-\mathbb{E}[\pi_{DP}^{blind}(X^{s},X^{\tau})\mid A=b]=0$. The fairness-value tradeoff is bounded by:
$V(\pi^{*})-V(\pi_{DP}^{blind})\leq\left|s(X,A)-s(X^{s},X^{\tau})\right|\leq\frac{(1-\bar{p}_{0})\left|\theta^{\kappa}\right|\mathbb{E}[\left|\tau(X^{s},X^{\tau})\right|]\mathbb{E}[\left|(X^{\kappa}-\mathbb{E}[X^{\kappa}])\right|}{\sqrt{2\pi}}.$
with the last inequality holding in the above model of Equation 5.
Remark 9.2
Proposition 9.1 characterizes fairness-value trade-offs in the simple proxy compliance model. The disparity $\Delta_{\kappa}=\left|\mathbb{E}[\kappa(X,A)\mid A=1]-\mathbb{E}[\kappa(X,A)\mid A=0]\right|$ captures the compliance disparity induced by $X^{\kappa}$, which we interpret as a proxy for the fairness benefit of dropping $X^{\kappa}$. (It is a proxy because the fairness benefit depends on the budget in general). In contrast, $\theta^{\kappa}\mathbb{E}|X^{\kappa}-\mathbb{E}[X^{\kappa}]|.$ Proxy-blindness achieves low fairness-welfare trade-off when $X^{\kappa}$ is mostly a marker of group difference, rather than mostly a source of individual-level compliance prediction within each group.
Interpretation
We set the parameter vector to $\beta_{\tau}=1.25$, $\lambda^{s}=0.9$ (treatment), $\mu_{a}=2$, $\mu_{b}=-2$, $\lambda^{\kappa}=0.18$, $\bar{p}_{0}=0.2$, and $c_{\kappa}=0.5$. The unconstrained policy $\pi^{*}$ attains value $V(\pi^{*})=0.2159$ at budget $b=0.3$. Under the unconstrained policy, encouragement disparity is $\mathbb{E}[\pi^{*}(X,A)\mid A=a]-\mathbb{E}[\pi^{*}(X,A)\mid A=b]=0.1232$ and treatment disparity is $|\mathbb{E}[T(\pi^{*})\mid A{=}a]-\mathbb{E}[T(\pi^{*})\mid A{=}b]|=0.0652$. A compliance-proxy blind policy $\pi^{\mathrm{blind}}$ (ranking by $(X^{s},X^{\tau})$ alone) attains $V(\pi^{\mathrm{blind}})=0.2128$ with encouragement disparity $0.0088$ and treatment disparity $0.0328$. Therefore the value gap from blinding is $V(\pi^{*})-V(\pi^{\mathrm{blind}})=0.0031$ (a relative 1.4% drop) although it eliminates the $0.1232$ encouragement disparity. Therefore, in this simple model, covariate-rich personalized policies offer more degrees of freedom to find decision alternatives that could greatly reduce outreach disparities at modest loss of value. Of course, the extent of such trade-offs in general depends on the data and context in any particular application.
Table 2: Budget–access–value tradeoffs at budget $b=0.30$. Compliance disparity $\Delta_{\kappa}=0.1071$.
Policy $V(\cdot)$ $V(\pi^{*}){-}V(\cdot)$ Budget disp. Treatment disp.
$\pi^{*}$ 0.2159 0 0.1232 0.0652
$\pi^{blind}$ 0.2128 0.0031 0.0088 0.0328
10 Additional discussion on method
10.1 Extrapolation in 3.4
@@ 10 Additional discussion on method
Proposition 10.1 (Control variate for alternative identification )
@@
$\textstyle V(\pi)=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}\left[\left\{Y(t)\frac{\mathbb{I}\left[T=t\right]}{p_{t}(X)}+\left(1-\frac{\mathbb{I}\left[T=t\right]}{p_{t}(X)}\right)\mu_{t}(X)\right\}p_{t\mid r}(X)\right]$
@@
11 Additional discussion on general optimization method
11.1 Additional discussion on constrained optimization
@@
$\max_{\pi}\;\{V(\pi)\colon\mathbb{E}[T(\pi)\mid A=a]-\mathbb{E}[T(\pi)\mid A=b]\leq\epsilon\}$ (8)
11.2 Additional discussion on Algorithm 1 (general algorithm)
11.2.1 Constrained policy optimization oracle
Although Agarwal et al. (2018) use the more general notation $Mh(\pi)\leq d$, in our setting these constraints are encoded directly as coordinates of $\Delta(\pi)$, so we use $\Delta(\pi)\leq d$ in the main text. Agarwal et al. [2018] consider a generic formulation of fairness constraints as moment conditions on $\pi$. Their key algorithmic design choices include convexification of the space of policies, that is, they optimize over distributions over policies $Q\in\mathcal{P}(\Pi),$ where $\mathcal{P}(\Pi)$ is the space of distributions over policies. The linear constraints are encoded as coordinates $\Delta_{k}(\pi)\leq d_{k}$ of the moment vector $\Delta(\pi)$, with observed-data scores $\delta_{k}(O;\pi,\eta)$ satisfying $\mathbb{E}[\delta_{k}(O;\pi,\eta_{0})]=\Delta_{k}(\pi)$. The elements of $\Delta_{k}(\pi)$ are average functionals (for example, the average treatment takeup in group $k$). Importantly, the moment function $\delta_{k}$ depends on $\pi$, while the conditioning event cannot depend on $\pi$. Many important fairness constraints can nonetheless be written in this framework, such as burden/resource parity and parity in true positive rates, but not measures such as calibration whose conditioning event does depend on $\pi$. (See Section 11.2 for examples omitted for brevity.)
Implementing best-response oracle
Given $\lambda_{t},$ the algorithm computes a best response over $Q$ ($\operatorname{BEST}_{\beta}\left({\lambda}_{t}\right)$); since the worst-case distribution will place all its weight on one classifier, this step reduces to cost-sensitive/weighted classification [Beygelzimer and Langford, 2009, Zhao et al., 2012], which we describe in further detail below. Computing the best response over $\operatorname{BEST}_{{\lambda}}(\hat{Q}_{t}))$ selects the most violated constraint.
11.2.2 Weighted classification reduction and off-policy estimation
@@
11.2.3 Additional fairness constraints and examples in this framework
In this section we discuss additional fairness constraints and how to formulate them in the generic framework. Much of this discussion is quite similar to [Agarwal et al., 2018] (including in notation) and is included in this appendix for completeness only. We additionally provide concrete discussion of the reduction to weighted classification, and concrete descriptions of the causal fairness constraints in the more general framework.
@@ 11.2.3 Additional fairness constraints and examples in this framework
Example 11.1 (Writing treatment parity in the general constrained classification framework.)
@@
The score $\delta_{1}(O;\pi,\eta)$ that identifies the upper constraint $\Delta_{1}(\pi)=\mathbb{E}[T(\pi)\mid A=a]-\mathbb{E}[T(\pi)\mid A=b]$ is
$\delta_{1}(O;\pi,\eta)=\frac{\mathbf{1}\{A=a\}}{p_{a}}\sum_{r}\pi_{r}(X)\,p_{1\mid r}(X,A)-\frac{\mathbf{1}\{A=b\}}{p_{b}}\sum_{r}\pi_{r}(X)\,p_{1\mid r}(X,A),$
where $p_{u}:=P(A=u)$, and $\delta_{2}(O;\pi,\eta)=-\delta_{1}(O;\pi,\eta)$ for the lower constraint $\Delta_{2}(\pi)\leq\varepsilon$. One can verify $\mathbb{E}[\delta_{k}(O;\pi,\eta_{0})]=\Delta_{k}(\pi)$ coordinatewise.
Both scores are linear in the policy: using $\sum_{r}\pi_{r}(X)\,p_{1\mid r}(X,A)=\pi_{1}(X)\,(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))+p_{1\mid 0}(X,A)$, we can write $\delta_{1}(O;\pi,\eta)=\pi_{1}(X)\,\tilde{w}(O;\eta)+c(O;\eta)$ for known weight $\tilde{w}$ and offset $c$ that depend only on nuisances, not on $\pi$. This linearity in $\pi$ implies linearity of $\Delta_{k}(Q)=\mathbb{E}[\delta_{k}(O;Q,\eta_{0})]$ in the randomized policy $Q$, which is what enables the saddle-point reductions approach of Agarwal et al. [2018]. This yields two coordinates of $\Delta(\pi)$, corresponding to the upper and lower inequality constraints.
In binary monotone settings, the same framework can also encode responder-conditional constraints, but we omit the details because they are not used in the main text or case studies.
11.2.4 Best-response oracles